19. A box weighing 15 lb is pulled by a force of 10 lb along a plane inclined 30o with the horizontal, the force being parallel to the plane. Starting from rest, how far will the box travel after 5 sec? 


1
Expert's answer
2021-09-28T11:14:36-0400

Write the equations by Newton's second law by axes:


"ma=-W\sin\theta+F,\\\space
a=\frac{F-W\sin\theta}{m}."

The distance is


"d=\frac{at^2}{2}=\frac{(F-W\sin\theta)t^2}{2m}=67\text{ ft}."